OpenStudy (anonymous):

This problem is having me find a value for y on the graph? The equation is y = x^2 + 2x -2 with -3 for x? How do you solve for y?

OpenStudy (anonymous):

you plug in -3 for x

OpenStudy (anonymous):

plug x = -3 in the eqn

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so it would be y=(-3)^2 +2(-3) -2 y=9-6-2 y=1